OpenStudy (anonymous):

7x + 15/3-2 = 3 ( x-2) + 7 +2x

OpenStudy (lilg132):

first you open the brackets 7x + 15/3 - 2 = 3(x - 2) + 7 + 2x 7x + 15/3 - 2 = 3x - 6 + 7 + 2x now we need to bring the letters to one side and numbers to the other 7x - 3x - 2x = 7 - 6 - 15/3 + 2 which gives us 2x = 2.2 to get the value of x we need to get x on its own so we divide it by the number next to it and if we divide one side we also divide the other side 2x/2 = 2.2 / 2 gives you x = 1.1 Dont forget to click best response =)